# Modelling Quantum Transduction for Multipartite Entanglement Distribution

### v1

@@ III-A Quantum Transduction Efficiency
Here we focus on electro-optical quantum transducers due to their attractive features – highlighted in Footnote 2 – but the theoretical analysis we develop can be easily extended to different transduction hardware, by properly accounting for the particulars in the expression of the conversion efficiency [^4] See as instance eq. 6 in [9] for opto-electro-mechanics transducers..
@@
Similarly, the efficiency for down-converting optical photons into microwave photos is denoted with $\eta_{\downarrow}$. In the following, we reasonably assume the up- and down-conversion efficiencies being symmetric [12, 27], i.e.:
@@
Regarding extraction ratios $\zeta_{x}$, values of $\zeta_{x}$ close to $1$ map into internal cavity losses $\gamma_{x,i}\stackrel{{\scriptstyle\triangle}}{{=}}\gamma_{x}-\gamma_{x,e}$ being relatively small [27]. It mush be noted that there is a wide-scientific consensus in considering unitary values for $\zeta_{x}$ feasible to achieve in the near-future. And indeed, typical values assumed in theoretical studies are around $\zeta_{x}=0.9$ [21], whereas experimental values in the order of $0.1-0.2$ have already been measured [34]. On the contrary, unitary cooperativity is still considered beyond the state-of-the art in the near-future, with experimental values for $C$ measured in the order of $10^{-3}-10^{-5}$ [12, 34]. In the following subsections, we study and discuss the impact of these hardware constraints on the key functionality represented by entanglement distribution.
@@ III-B TMD with “vanilla” quantum transduction
As mentioned at the beginning of this section, in TMD with vanilla quantum transduction [^6] In the following, when there is no ambiguity issue, we could refer to TMD strategy with vanilla quantum transduction as vanilla TMD., the ebit of each EPR – locally generated at the orchestrator – is distributed to the clients with a sequence of up- and down-conversion. Thus, we can model the TMD strategy with vanilla quantum transduction as in Fig. 2. Specifically, $n$ microwave EPR pairs |$\ket{\Phi^{o,o}_{M,M}}$ are generated at the orchestrator (with superscript |$\ket{\cdot^{o}}$ denoting the orchestrator and subscript |$\ket{\cdot_{M}}$ denoting the microwave photon domain), with one ebit of each EPR pair up-converted to optical domain, transmitted to each client and down-converted to the microwave domain therein, eventually resulting in an EPR state, denoted as |$\ket{\Phi^{o,c}_{M,M}}$, since it is shared between the orchestrator $o$ and the arbitrary client $c\in\{1,\ldots,n\}$. Accordingly, once all the $n$ EPRs are distributed through the network, the overall state |$\ket{\Omega^{v}}$ obtained with vanilla TMD before teleportation is given by:
@@
$p^{v}_{c}=\eta^{o}_{\uparrow}\eta^{c}_{\downarrow}e^{-\frac{l_{o,c}}{L_{o}}}$ (7)
with $l_{o,c}$ denoting the length of the fiber link between orchestrator and client $c$, $L_{o}$ denoting the attenuation length of the fiber [^7] As for today, commercial fibers feature an attenuation lower than $1$db/km. As instance, optical photons with wavelength equal to $1550$nm – i.e., DWDM ITU $100$GHz channel number $35$ in the C band – experience an attenuation of $0.2$dB/km, which corresponds to $L_{o}=22$km [36]. and $\eta^{o}_{\uparrow}$ and $\eta^{c}_{\downarrow}$ denoting the efficiency of the up- and down-conversion at the orchestrator and at the client, respectively.
@@
In Fig. 4, we plot the ebit distribution probability $p^{v}_{c}$ as a function of the cooperativity $C$ and the length $l_{o,c}$ of the fiber between orchestrator and arbitrary client $c$ for both DMD and vanilla-TMD. For computing the probability, we have reasonably assumed similar transduction hardware for both up- and down-conversion, since the two conversion processes present similar hardware challenges. Specifically, we have assumed symmetric efficiencies – i.e., $\eta^{c}_{\downarrow}=\eta^{o}_{\uparrow}$ – characterized by identical cooperativity, i.e., $C=C_{\uparrow}=C_{\downarrow}$. Furthermore, we have considered ideal extraction ratios, i.e., $\zeta_{o}=\zeta_{m}=1$. This assumption is not restrictive since, as pointed out in the previous subsection, it is commonly adopted in literature and experimental values with almost similar order of magnitude have been measured. Clearly, the results in Fig. 4 can be easily adapted to non-ideal extraction ratio by simply weighting the probability value with the appropriate scaling factor $(\zeta_{o}\zeta_{m})^{2}$.
@@
It may seem that the overall transduction process underlying vanilla-TMD on a single orchestrator-client link – i.e., a cascade of up- and down-conversion – coincides with the overall process of DMD on a single orchestrator-client link. However, the fundamental difference is that the former acts on the ebit of an EPR pair whereas the latter acts on the ebit of the multipartite entangled state. This implies that the probability given in (7) can be used as-is for modelling the probability of successful distribution of a single ebit via DMD. However, such a probability does not capture the persistency properties of the considered multipatite entanglement class. In other words, in DMD the successful distribution of an ebit on a certain orchestrator-client link does not assure the preservation of the overall multipartite entangled state, and thus it does not assure the success of the overall transduction process. As a consequence, the successful distribution of an ebit on a certain orchestrator-client link is only a necessary condition for the success the overall transduction process in DMD. In this light, in Lemma 2, we provide the necessary condition for having in DMD a non-null quantum capacity on a given orchestrator-client link.
@@
From Lemma 2, it results that for having unitary capacity on a given orchestrator-client link, $p^{v}_{c}$ should be unitary as well. By assuming fiber length significantly shorter than the attenuation length $L_{0}$, this requires unitary cooperativity $C=1$ for both the up- and the down- transducers. Yet, such a value, as pointed out in Sec. III-A, exceeds current state-of-the-art technologies of several orders of magnitudes. Hence, “noiseless” DMD is beyond current quantum transduction capabilities, regardless of the length of the fiber connecting orchestrator with client.
One might believe that at least “noisy”DMD – i.e., DMD characterised by non-null capacity on a given orchestrator-client link – should be feasible, by restricting the multipartite entangled state to classes characterized by high persistence. Unfortunately this is not true since, for satisfying the condition $p^{v}_{c}>\frac{1}{2}$, the cooperativity values – by combining the effects of both up- and down-conversion – should satisfy $C>2\sqrt{2}-2\sqrt{2-\sqrt{2}}-1\simeq 0.30$, which are again beyond current quantum transduction capabilities. In Fig. 4, we denoted the values for the cooperatives $C=C_{\uparrow}=C_{\downarrow}$ enabling non-null one-way quantum capacity with the dotted black curve.
@@
The result of Lemma 3 induces us to discuss some key aspects. First, the condition in Lemma 3 characterizes the entire distribution of the multipartite state, differently from Lemma 2, for the reasons highlighted above. Second, given that the two-way quantum capacity $Q_{2}$ – which, as result of Lemma 3. is equal to $Q_{2}=p^{v}_{c}$ – coincides with both the one-way and the two-way entanglement distribution capacity $D_{1}$ and $D_{2}$ [39, 40], it results that an average of $p^{v}_{c}>0$ ebits can be distilled from each distribution attempt, or equivalently $\frac{1}{p^{v}_{c}}$ distribution attempts must be performed in average to distil one EPR. Clearly, as highlighted for DMD, in these consideration we are not accounting for the effects of the encoder implemented within the transducer. If these effects should be accounted for, then $p^{v}_{c}$ represents an upper bound for quantifying the distillable entanglement. From Fig.4, it is evident that vanilla TMD does not require unitary cooperativity for achieving unitary capacities. This is remarkable with respect to DMD.
@@ III-C TMD with intrinsic-entanglement quantum transduction (IE-TMD)
Conversely, the parameter regime achievable with state-of-the-art technology – when coupled with cryogenic temperatures so that thermal microwave noise can be neglected [27] – enables the generation of intrinsic entanglement, i.e., entanglement between two different (optical and microwave) domains [21, 27], as shown in Fig. 5. Specifically, through spontaneous parametric down-conversion  [^8] Spontaneous parametric down-conversion is a non-linear optical process where a photon spontaneously splits into two photons of lower energies [42]. (SPDC) of an input pump field, entanglement between optical and microwave fields is generated within the transducer [21, 28, 29, 43]
@@
|$\ket{\Phi^{o,o}_{M,O}}=\frac{1}{\sqrt{2}}(\ket{0_{M}^{o}0_{O}^{o}}+\ket{1_{M}^{o}1_{O}^{o}}).$ (8)
with the subscripts $(\cdot_{M})$ and $(\cdot_{O})$ denoting the photon domain – i.e., microwave or optical – and the superscripts $(\cdot^{o})$ and $(\cdot^{c})$ denoting the “location” of each ebit, i.e., at the orchestrator or at the client. Accordingly, in (8) the term |$\ket{1_{M}^{o}1_{O}^{o}}$ denotes the event in which SPDC successfully generated a couple of photons, one at microwave and the other at optical frequency. Conversely, the term |$\ket{0_{M}^{o}0_{O}^{o}}$ denotes the event in which SPDC failed, and no microwave nor optical photon have been generated. Entanglement in the form of (8) is generated whenever the quantum transducer is initialized with no input – namely, no photon to be converted as in Fig. 5 – and the input pump field is set to operate on a frequency that is the sum of the frequencies of the optical and microwave photons (aka “blue detuning”).
@@
Two assumptions underlying equation (8) must be better discussed. Regarding the assumption of generating a two-level state – i.e., the assumption of restricting the admissible system state to 2-level Fock states for each field rather than an uncountable levels – it is not restrictive, since it simply requires a specific inizialization (aka “red detuning”) of the microwave field inside the cavity as in Fig. 5 [43]. This leads to an EPR state in the form of |$\ket{\Psi^{o,o}_{M,O}}=\frac{1}{\sqrt{2}}(\ket{0^{o}_{M}1^{o}_{O}}+\ket{1^{o}_{M}0^{o}_{O}})$, which is equivalent to (8) up to a basis change. Thus, in the following we will use (8) for the sake of simplicity. Furthermore, also the assumption of an EPR state – i.e., a maximally entangled state with even superposition of two states as in (8) – depends on a careful setting of the transduction hardware parameters [43]. Obviously, any hardware mismatch from the ideal setting would impact on the purity of the generated entangled pair. And we will properly account for this mismatch with the following Lemma 4.
@@
Accordingly, once all the $n$ EPRs are distributed through the network, the overall state |$\ket{\Omega^{IE}}$ obtained with IE-TMD is equivalent to the state given in (6), and the teleportation of the multipartite entangled state can be now be performed.
@@
$p^{IE}_{c}=S\big{(}\eta^{o}_{\uparrow}\big{)}\eta^{c}_{\downarrow}e^{-\frac{l_{o,c}}{L_{o}}}$ (10)
where $l_{o,c}$ denotes the length of the fiber link between orchestrator and client $c$, $L_{o}$ denotes the attenuation length of the fiber, $\eta^{o}_{\uparrow}$ denotes the efficiency of the intrinsic entanglement transducer at the orchestrator, $\eta^{c}_{\downarrow}$ denotes the efficiency of the down-conversion at the client, and $S(\cdot)$ [^9] With a small abuse of notation, we have indicated in the argument of the Von Neuman entropy the eigenvalue determining its value rather than – as usually done – the density matrix on which the entropy is evaluated. denotes Von Neuman entropy given by:
@@
Before discussing the implication of Lemma 4, as highlighted in the appendix it is worthwhile to clarify that – although the physical mechanisms are different – $\eta^{o}_{\uparrow}$ in (3) still governs the efficiency of the intrinsic transducer at the orchestrator, since the hardware is the same. But the different physical mechanics imply less stringent requirements on the cooperativity, as discussed in the following.
In Fig. 7, we plot the successful ebit distribution probability $p^{IE}_{c}$ given in (11) as a function of the cooperativity $C$ and the length $l_{o,c}$ of the fiber between orchestrator and the arbitrary client $c$. As mentioned above, for computing the probability, we have reasonably assumed similar transduction hardware for both up- and down-conversion, since the two conversion processes present similar hardware challenges. Yet, as discussed in the proof of Lemma 4, there exists a subtle difference in the desired operative parameters between the transducer at the orchestrator and that one at the client. Specifically, while at the client we target a unitary conversion efficiency $\eta^{c}_{\downarrow}$, at the orchestrator we aim to obtain a conversion efficiency $\eta^{o}_{\uparrow}$ equal to $\frac{1}{2}$ – which corresponds to $C=C_{th}=3-2\sqrt{2}$ – so that an EPR pair can be generated. From (3), we have that $\eta^{o}_{\uparrow}$ monotonically increases with cooperativity $C$, as visually confirmed by Fig. 5. However, as discussed in Sec III-A, the challenge is toward obtaining higher values of $C$ rather than lowers. Hence, we avoid foolish parameter setting by computing $\eta^{o}_{\uparrow}$ as a function of $\min\{C,C_{th}\}$. Clearly, this choice – reasonable from a communication perspective aiming at optimizing the entanglement distribution process – maps into having orchestrator and client characterized by different hardware parameter setting. In this context, it is worth noting that – differently from what would be desirable – transducers at clients requires higher (hence, harder to achieve) cooperativity values than transducers at the orchestrator. Furthermore, in agreement with Sec. III-B, we have considered ideal extraction ratios, i.e., $\zeta_{o}=\zeta_{m}=1$.

### v2

@@ III-A Quantum Transduction Efficiency
Here we focus on electro-optical quantum transducers due to their attractive features – highlighted in Footnote 3 – but the theoretical analysis we develop can be easily extended to different transduction hardware, by properly accounting for the particulars in the expression of the conversion efficiency [^6] See as instance eq. 6 in [9] for opto-electro-mechanics transducers..
@@
Similarly, the efficiency for down-converting optical photons into microwave photos is denoted with $\eta_{\downarrow}$. In the following, we reasonably assume the up- and down-conversion efficiencies being symmetric [12, 33, 47], i.e.:
@@
Regarding extraction ratios $\zeta_{x}$, values of $\zeta_{x}$ close to $1$ map into internal cavity losses $\gamma_{x,i}\stackrel{{\scriptstyle\triangle}}{{=}}\gamma_{x}-\gamma_{x,e}$ being relatively small [33]. It mush be noted that there is a wide-scientific consensus in considering unitary values for $\zeta_{x}$ feasible to achieve in the near-future. And indeed, typical values assumed in theoretical studies are around $\zeta_{x}=0.9$ [32], whereas experimental values in the order of $0.1-0.2$ have already been measured [48]. On the contrary, unitary cooperativity is still considered beyond the state-of-the art in the near-future, with experimental values for $C$ measured recently reaching around $0.3$ [34]. In the following subsections, we study and discuss the impact of these hardware constraints on the key functionality represented by entanglement distribution.
Remark.
It is evident from the previous equations, that quantum transduction is not just a merely frequency conversion process. Many factors concur to challenge the interfacing between different hardware platforms like photonic and superconducting. For instance, in order to achieve a good transduction, the physical modes of microwave and optical systems must be matched, which includes considerations of impedance, spatial overlap, and the temporal properties of the signals. This, as showed in [11], can be captured by the electro-optic coupling coefficient $g$ in (1). Accordingly, in the following we focus on the conversion efficiency in order to abstract from the hardware particulars. And in Sec. IV, we provide a more detailed discussion about this aspect.
@@ III-B TMD with “vanilla” quantum transduction
As mentioned at the beginning of this section, in TMD with vanilla quantum transduction [^8] In the following, when there is no ambiguity issue, we could refer to TMD strategy with vanilla quantum transduction as vanilla TMD., one ebit of each EPR – locally generated at the orchestrator – is distributed to the clients with a sequence of up- and down-conversion. Thus, we can model the TMD strategy with vanilla quantum transduction as in Fig. 2. Specifically, $n$ microwave EPR pairs |$\ket{\Phi^{o,o}_{M,M}}$ are generated at the orchestrator, where the superscript |$\ket{\cdot^{o}}$ denotes the orchestrator as “location” of each ebit of the EPR and the subscript |$\ket{\cdot_{M}}$ denotes the microwave photon domain of each ebit. Once the generation is performed, one ebit of each EPR pair generated is up-converted to optical domain, transmitted to each client and down-converted to the microwave domain therein. The process eventually results in an EPR state, denoted as |$\ket{\Phi^{o,c}_{M,M}}$, since it is shared between the orchestrator $o$ and the arbitrary client $c\in\{1,\ldots,n\}$. Accordingly, once all the $n$ EPRs are distributed through the network, the overall state |$\ket{\Omega^{v}}$ obtained with vanilla TMD before teleportation is given by:
@@
$p^{v}_{c}=\eta^{o}_{\uparrow}\eta^{c}_{\downarrow}e^{-\frac{l_{o,c}}{L_{0}}}$ (7)
with $l_{o,c}$ denoting the length of the fiber link between orchestrator and client $c$, $L_{0}$ denoting the attenuation length of the fiber [^9] As for today, commercial fibers feature an attenuation lower than $1$db/km. As instance, optical photons with wavelength equal to $1550$nm – i.e., DWDM ITU $100$GHz channel number $35$ in the C band – experience an attenuation of $0.2$dB/km, which corresponds to $L_{0}=22$ km [50]. and $\eta^{o}_{\uparrow}$ and $\eta^{c}_{\downarrow}$ denoting the efficiency of the up- and down-conversion at the orchestrator and at the client, respectively.
@@
In (5), we have reasonably assumed that the processes of up and down conversion are symmetrical. However, it is worth to highlight that the quality of the optical ebit before the down-conversion process could be smaller than the quality of microwave ebits before up-conversion, due to the degradation effects introduced by the optical communication channel. Therefore, in Lemma 1, we modelled the decay – exponential with the distance – experienced by an entangled photon travelling over an optical communication channel with the term $e^{-\frac{l_{o,c}}{L_{0}}}$. This allows us to capture the more challenging nature of the down conversion process.
In Fig. 4, we plot the ebit distribution probability $p^{v}_{c}$ as a function of the cooperativity $C$ and the length $l_{o,c}$ of the fiber between orchestrator and arbitrary client $c$ for both DMD and vanilla-TMD. For computing the probability, we have reasonably assumed similar transduction hardware for both up- and down-conversion, since the two conversion processes present similar hardware challenges. Specifically, we have assumed symmetric efficiencies – i.e., $\eta^{c}_{\downarrow}=\eta^{o}_{\uparrow}$ – characterized by identical cooperativity, i.e., $C=C_{\uparrow}=C_{\downarrow}$. Furthermore, we have considered ideal extraction ratios, i.e., $\zeta_{o}=\zeta_{m}=1$. This assumption is not restrictive since, as pointed out in the previous subsection, it is commonly adopted in literature and experimental values with almost similar order of magnitude have been measured. Clearly, the results in Fig. 4 can be easily adapted to non-ideal extraction ratio by simply weighting the probability value with the appropriate scaling factor $(\zeta_{o}\zeta_{m})^{2}$. It may seem that the overall transduction process underlying vanilla-TMD on a single orchestrator-client link – i.e., a cascade of up- and down-conversion – coincides with the overall process of DMD on a single orchestrator-client link. However, the fundamental difference is that the former acts on the ebit of an EPR pair whereas the latter acts on the ebit of the multipartite entangled state. This implies that the probability given in (7) can be used as-is for modelling the probability of successful distribution of one EPR pair via TMD and of a single ebit of the multipartite state via DMD. However, in DMD such a probability does not capture the persistency properties of the considered multipatite entanglement class. In other words, in DMD the successful distribution of an ebit on a certain orchestrator-client link does not assure the preservation of the overall multipartite entangled state, and thus it does not assure the success of the overall transduction process. As a consequence, the successful distribution of an ebit on a certain orchestrator-client link is only a necessary condition for the success the overall transduction process in DMD. In this light, in Lemma 2, we provide the necessary condition for having in DMD a non-null quantum capacity on a given orchestrator-client link.
@@
From Lemma 2, it results that for having unitary capacity on a given orchestrator-client link, $p^{v}_{c}$ should be unitary as well. By assuming fiber length significantly shorter than the attenuation length $L_{0}$, this requires unitary cooperativity $C=1$ for both the up- and the down- transducers. Yet, such a value, as pointed out in Sec. III-A, exceeds current state-of-the-art technologies.
Hence, “noiseless” DMD is beyond current quantum transduction capabilities, regardless of the length of the fiber connecting orchestrator with client.
One might believe that at least “noisy” DMD – i.e., DMD characterised by non-null capacity on a given orchestrator-client link – should be feasible, by restricting the multipartite entangled state to classes characterized by high persistence. Unfortunately this is not true since, for satisfying the condition $p^{v}_{c}>\frac{1}{2}$, the cooperativity values – by combining the effects of both up- and down-conversion – should satisfy $C>2\sqrt{2}-2\sqrt{2-\sqrt{2}}-1\simeq 0.30$.
In Fig. 4, we denoted the values for the cooperatives $C=C_{\uparrow}=C_{\downarrow}$ enabling non-null one-way quantum capacity with the dotted black curve.
@@
The result of Lemma 3 induces us to discuss some key aspects. First, the condition in Lemma 3 characterizes the entire distribution of the multipartite state, differently from Lemma 2, for the reasons highlighted above. Second, given that the two-way quantum capacity $Q_{2}$ – which, as result of Lemma 3. is equal to $Q_{2}=p^{v}_{c}$ – coincides with both the one-way and the two-way entanglement distribution capacity $D_{1}$ and $D_{2}$ [53, 54], it results that an average of $p^{v}_{c}>0$ ebits can be distilled from each distribution attempt, or equivalently $\frac{1}{p^{v}_{c}}$ distribution attempts must be performed in average to distil one EPR pair. Clearly, as highlighted for DMD, in these considerations we are not accounting for the effects of the encoder implemented within the transducer. If these effects were taken into account, then $p^{v}_{c}$ represents an upper bound for quantifying the distillable entanglement. From Fig. 4, it is evident that vanilla TMD does not require unitary cooperativity for achieving unitary capacities. This is remarkable with respect to DMD.
@@ III-C TMD with intrinsic-entanglement quantum transduction (IE-TMD)
Conversely, the parameter regime achievable with state-of-the-art technology – when coupled with cryogenic temperatures so that thermal microwave noise can be neglected [33] – enables the generation of intrinsic entanglement, i.e., entanglement between two different (optical and microwave) domains [32, 33], as shown in Fig. 5. Specifically, through spontaneous parametric down-conversion  [^10] Spontaneous parametric down-conversion is a non-linear optical process where a photon spontaneously splits into two photons of lower energies [56]. (SPDC) of an input pump field, entanglement between optical and microwave fields is generated within the transducer [32, 41, 42, 57].
@@
|$\ket{\Phi^{o,o}_{M,O}}=c_{0}\ket{0_{M}^{o}0_{O}^{o}}+c_{1}\ket{1_{M}^{o}1_{O}^{o}}.$ (8)
with the subscripts $(\cdot_{M})$ and $(\cdot_{O})$ denoting the photon domain – i.e., microwave or optical – and the superscripts $(\cdot^{o})$ and $(\cdot^{c})$ denoting the “location” of each ebit, i.e., at the orchestrator or at the client. Accordingly, in (8) the term |$\ket{1_{M}^{o}1_{O}^{o}}$ denotes the event in which SPDC successfully generated a couple of photons, one at microwave and the other at optical frequency. Conversely, the term |$\ket{0_{M}^{o}0_{O}^{o}}$ denotes the event in which SPDC failed, and no microwave nor optical photon have been generated. The coefficients $c_{0}$ and $c_{1}$ depends on the transducer hardware parameters. Entanglement in the form of (8) is generated whenever the quantum transducer is initialized with no input – namely, no photon to be converted as in Fig. 5 – and the input pump field is set to operate on a frequency that is the sum of the frequencies of the optical and microwave photons (aka “blue detuning”).
@@
Two assumptions underlying equation (8) must be better discussed. Regarding the assumption of generating a two-level state – i.e., the assumption of restricting the admissible system state to 2-level Fock states for each field rather than an uncountable levels – it is not restrictive, since it simply requires a specific inizialization of the microwave field inside the cavity as in Fig. 5 [57]. In this case, an input pump field is set to operate on a frequency that is the difference of the frequencies of the optical and microwave photons (aka “red detuning”). The pump is exploited to obtain an entangled state in the form of |$\ket{\Psi^{o,o}_{M,O}}=c_{0}\ket{0^{o}_{M}1^{o}_{O}}+c_{1}\ket{1^{o}_{M}0^{o}_{O}}$ , which is equivalent to (8) up to a basis change [57, 61].
The term |$\ket{0_{M}^{o}1_{O}^{o}}$ denotes a successful conversion of the initialized microwave field within the transducer, while the term |$\ket{1_{M}^{o}0_{O}^{o}}$ denotes a failed conversion. Again, the coefficients $c_{0}$ and $c_{1}$ depends on the transducer features. Depending on a careful setting of the transduction hardware parameters, it is possible to obtain an EPR state (see Appendix C) in the form:
|$\ket{\Psi^{o,o}_{M,O}}=\frac{1}{\sqrt{2}}(\ket{0^{o}_{M}1^{o}_{O}}+\ket{1^{o}_{M}0^{o}_{O}})$ (9)
Obviously, any hardware mismatch from the ideal setting would impact on the purity of the generated entangled pair. And we will properly account for this mismatch with the following Lemma 4. Thus, in the following we will use (9) for the sake of simplicity.
@@
Accordingly, once all the $n$ EPR pairs are distributed through the network, the overall state |$\ket{\Omega^{IE}}$ obtained with IE-TMD is equivalent to the state given in (6), and the teleportation of the multipartite entangled state can be now be performed.
@@
$p^{IE}_{c}=S\big{(}\eta^{o}_{\uparrow}\big{)}\eta^{c}_{\downarrow}e^{-\frac{l_{o,c}}{L_{0}}}$ (11)
where $l_{o,c}$ denotes the length of the fiber link between orchestrator and client $c$, $L_{0}$ denotes the attenuation length of the fiber, $\eta^{o}_{\uparrow}$ denotes the efficiency of the intrinsic entanglement transducer at the orchestrator, $\eta^{c}_{\downarrow}$ denotes the efficiency of the down-conversion at the client, and $S(\cdot)$ [^11] With a small abuse of notation, we have indicated in the argument of the Von Neuman entropy the eigenvalue determining its value rather than – as usually done – the density matrix on which the entropy is evaluated. denotes Von Neuman entropy given by:
@@
Before discussing the implication of Lemma 4, as highlighted in the appendix it is worthwhile to clarify that – although the physical mechanisms are different – $\eta^{o}_{\uparrow}$ in (3) still governs the efficiency of the intrinsic transducer at the orchestrator, since the hardware is the same. But the different physical mechanics imply less stringent requirements on the cooperativity, as discussed in the following. In Fig. 7, we plot the successful ebit distribution probability $p^{IE}_{c}$ given in (11) as a function of the cooperativity $C$ and the length $l_{o,c}$ of the fiber between orchestrator and the arbitrary client $c$. As mentioned above, for computing the probability, we have reasonably assumed similar transduction hardware for both up- and down-conversion, since the two conversion processes present similar hardware challenges. Yet, as discussed in the proof of Lemma 4, there exists a subtle difference in the desired operative parameters between the transducer at the orchestrator and that one at the client. Specifically, while at the client we target a unitary conversion efficiency $\eta^{c}_{\downarrow}$, at the orchestrator we aim to obtain a conversion efficiency $\eta^{o}_{\uparrow}$ equal to $\frac{1}{2}$ – which corresponds to $C=C_{th}=3-2\sqrt{2}$ – so that an EPR pair can be generated. From (3), we have that $\eta^{o}_{\uparrow}$ monotonically increases with cooperativity $C$, as visually confirmed by Fig. 3. However, as discussed in Sec III-A, the challenge is toward obtaining higher values of $C$ rather than lowers. Hence, we avoid foolish parameter setting by computing $\eta^{o}_{\uparrow}$ as a function of $\min\{C,C_{th}\}$. Clearly, this choice – reasonable from a communication perspective aiming at optimizing the entanglement distribution process – maps into having orchestrator and client characterized by different hardware parameter setting. In this context, it is worth noting that – differently from what would be desirable – transducers at clients requires higher (hence, harder to achieve) cooperativity values than transducers at the orchestrator. Furthermore, in agreement with Sec. III-B, we have considered ideal extraction ratios, i.e., $\zeta_{o}=\zeta_{m}=1$.